### v2

# Fixed-relative-switch strategies for learning based event-triggered control of nonlinear multiagent systems

Abstract
This paper investigates event-triggered control for consensus tracking in nonlinear semi-strict-feedback multi-agent systems (MASs) with unknown states and subject to disturbances. We begin by employing radial basis function neural networks combined with the backstepping method to approximate the unknown nonlinear dynamics of the MASs, which facilitates the development of state and disturbance observers for estimation. We then propose three adaptive event-triggered control strategies: fixed-threshold, relative-threshold, and switch-threshold, whose controllers and triggering conditions are defined acocrdingly. By applying Lyapunov stability theory, we rigorously prove that all follower agents’ outputs consistently track the reference of their leader, and all error signals remain uniformly bounded. Moreover, our control strategies effectively mitigate the occurrence of Zeno behaviors. To validate our control framework, an illustrative example is provided to demonstrate its desirable performance in consensus tracking and triggering efficiency.
Index Terms:
Multiagent system, adaptive control, observer, event-triggered control.
I Introduction
For decades, consensus tracking in nonlinear multi-agent systems (MASs) has been intensively studied in control and learning communities due to its wide applications in fields such as power grids [1][2], intelligent transportation [3][4] and communication systems [5][6]. The primary objective is to ensure that all units within the network operate ”unanimously” along a desired trajectory. However, uncertain system parameters and external perturbations in real engineering scenarios pose significant challenges for the problem and existing methods often fail to work. Consequently, it is critical to develop control strategies that effectively manage unknown states and perturbations.
In the conventional setting of sample-data control system, controllers continuously respond to changes in the system, which often results in unnecessary resource usage, particularly when the network has restricted bandwidth. A pivotal study in 1999 [9] demonstrated the advantages of event-triggered control over periodic impulse control in first-order stochastic systems with multiplicative noise, leading to its widespread adoption in both linear and nonlinear systems. Recent innovative design methodologies, as evidenced by works such as [7]-[16] have significantly advanced the field of event-triggered control.
Event-triggered control was initially applied to nonlinear systems and achieved great success in reducing trigger frequency and enhancing control performance, see, .e.g, [7] and [8]. The framework has then been extended in various directions, where new trigger conditions are introduced. For instance, [7] proposed a simple dynamic triggering condition that alleviates the traditional periodic execution requirements in closed-loop nonlinear systems. And [8] developed a technique that utilizes the current state of the plant to determine triggering time instances, without the need of periodicity. Also event-triggered mechanism has been extensively investigated in the context of MASs more recently. Specifically, [10] introduced a saturated-threshold event-triggered control strategy aimed at achieving robust consensus tracking in continuous-time nonlinear MASs under sensor attacks. And [11] employed a switching-based trigger strategy for multi-agent consensus tracking. Furthermore, [16] analyzed three event-triggered control strategies in nonlinear systems with uncertainty and did not require input-to-state stability (ISS).
Motivated by the above information, we have developed an adaptive consensus tracking framework for high-order nonlinear MASs. Our approach utilizes the backstepping method, filter method, RBF NNs, and observer method. It is specifically designed for scenarios where the states and perturbations are unknown, while only the output signals of the agents are available. The key contributions of this work are outlined below:
  • 1)
    We employ radial basis function neural networks and backstepping method to approximate the unknown transition functions of the MASs under investigation
  • 2)
    Based on the learned model, we develop both state and perturbation observers to monitor unknown states and unmeasured disturbances, which later enhance the robustness of the controller against perturbations.
  • 3)
    We propose three event-triggered control strategies: fixed-threshold, relative-threshold, and switch-threshold, together with a comprehensive analysis of their advantages such as reduced frequency of control updates, enhanced resource utilization and extended controller lifespan.
The organization of this paper is organized in the following way. Section II depicts the system model. The controller and event-triggered strategies are developed in Section III, while Section IV recommended the stability analysis. In Section V, we present our simulation results. Finally, we conclude the paper in Section VI.
II Preliminaries
II-A Problem Formulation
Consider a scenario with $M$ followers, designated as agents $1$ to $M$, and a leader, designated as agent 00, all interacting under a directed communication graph topology. Consider the following class of MASs [19][20]:
$\dot{x}_{i,k}$ $=x_{i,k+1}+f_{i,k}(\bar{x}_{i,k})+\xi_{i,k}$
$\dot{x}_{i,n}$ $=u_{i}+f_{i,n}(\bar{x}_{i,n})+\xi_{i,n}$
$y_{i}$ $=x_{i,1}$ (1)
where $i=1,...,N$, $j=1,...,n$, $k=1,...,n-1$, $\bar{x}_{i,j}=[x_{i,1},...,x_{i,j}]^{T}$ are the state of the $i$th follower, while $u_{i}\in R$ denotes its input of control. The output of the $i$th follower is expressed as $y_{i}\in R$. The function $f_{i,j}(\bar{x}_{i,j})$ represents $C^{1}$ class nonlinear smooth equation vectors. $\xi_{i,j}$ represents unmeasured external perturbations affecting the system. It is noted that The leader’s movement occurs independently, without being influenced by the actions or positions of the followers.
One intention of the works in this paper is to enable each follower’s output signal, ${{y}_{i}}$, to synchronously track the reference signal, ${{y}_{r}}$, through a suitable ETC strategy.
II-B Observer Design
In this paper, to effectively spot the unknown states and perturbations throughout the entire control system, we introduce a series of observers. By precisely analyzing and processing feedback data, these observers can accurately identify and estimate unknown variables within the system, thereby enhancing the overall robustness and stability.
The state observer is defined as
$\dot{\hat{\bar{x}}}_{i,n}$ $=(P_{i}\otimes I_{m})\dot{\hat{\bar{x}}}_{i,n}+(Q_{i}\otimes y_{i})+\sum_{l=1}^{n}(R_{i,l}\otimes\hat{f}_{i,l}(\dot{\hat{\bar{x}}}_{i,l}))$
$+(U_{i}\otimes u_{i})+\hat{\varpi}_{i}$
$y_{i}$ $=(V_{i}^{T}\otimes I_{m})\hat{\bar{x}}_{n}$ (2)
where $\otimes$ represents the Kronecker product, ${{\hat{\bar{x}}}_{i,l}}={{\left[\hat{x}_{i,1}^{T},\ldots,\hat{x}_{i,l}^{T}\right]}^{T}}$ represents the estimated value of the actual state. The vector $Q_{i}=[q_{i,1},...,q_{i,n}]^{T}$ is such that the matrix $P_{i}$ is a strictly Hurwitz. The parameters are defined as follows: $R_{i,1}=\left[\underbrace{0,\ldots,1}_{l},\ldots,0\right]_{n\times 1}^{T}$, $U_{i}=[0,...,0,1]^{T}_{n\times 1}$, $V_{i}=[1,0,...,0]^{T}_{n\times 1}$, $P_{i}=\left[\begin{gathered}-{q_{i,1}}\hfill\\ {\text{ }}\vdots{\text{ }}{I_{(n-1)\times(n-1)}}\hfill\\ -{q_{i,n}}{\text{ 0 }}\ldots{\text{ 0 }}\hfill\\ \end{gathered}\right]$.
Define $\psi_{i}=[(x_{i,1}-\hat{x}_{i,1})^{T},...,(x_{i,n}-\hat{x}_{i,n})^{T}]^{T}$ as the error of state observation, one has that
$\dot{\psi}_{i}=(P_{i}\otimes I_{m})\psi_{i}+\widetilde{\varpi}_{i}+\sum_{l=1}^{n}S_{i,l}\otimes(f_{i,l}(\overline{x}_{i,l})-\hat{f}_{i,l}(\hat{\overline{x}}_{i,l}))$ (3)
Define ${{\varphi}_{i}}(t)=\sum\limits_{l=1}^{n}{{{R}_{i,l}}\otimes({{f}_{i,l}}\left({{{\bar{x}}}_{i,l}}\left(t\right)\right)-{{{\hat{f}}}_{i,l}}({{{\hat{\bar{x}}}}_{i,l}}(t)))}$ as the error of function approximation and ${{\varphi}_{i}}(t)={{\left[\varphi_{i,1}^{T}(t),...,\varphi_{i,n}^{T}(t)\right]}^{T}}$, assume that ${{\varphi}_{i}}(t)$ is bounded, this implies that there exists an unspecified parameter $\varphi_{i}^{0}>0$ such that the inequality $\left\|{{\varphi_{i}}(t)}\right\|\leq\varphi_{i}^{0}$ holds true.
By utilizing RBF NNs to approximate the unidentified nonlinear function within the MASs (1), an optimal weight vector $W_{i,l}^{*}$ is defined as:
${{f}_{i,l}}({{\bar{x}}_{i,l}})=W_{i,l}^{*T}{{E}_{i,l}}\left({{{\hat{\bar{x}}}}_{i,l}}\right)+{{\sigma}_{i,l}}(t)$ (4)
where ${{\sigma}_{i,l}}(t)$ represents the bounded approximation error, i.e., there is a parameter ${{\sigma}_{0,l}}>0$ such that $\left|{{\sigma}_{i,l}}(t)\right|\leq{{\sigma}_{0,l}}$.
Consider $\hat{W}_{i,l}^{T}$ as the estimation of $W_{i,l}^{*T}$, the unidentified smooth nonlinear function can be estimated as
${{\hat{f}}_{i,l}}({{\hat{\bar{x}}}_{i,l}})=\hat{W}_{i,l}^{T}{E_{i,l}}\left({{{\hat{\bar{x}}}}_{i,l}}\right)$ (5)
The optimal weight $W_{i,l}^{*}$, with $l=1,\ldots,n$, is designed as follows
[…]
The state observers can be revised as follows
${{{\dot{\hat{x}}}}_{i,k}}={{{\hat{x}}}_{i,k+1}}+\hat{W}_{i,k}^{T}{{E}_{i,k}}\left({{{\hat{\bar{x}}}}_{i,k}}\right)+{{q}_{i,k}}{{\psi}_{i,1}}+{{{\hat{\varpi}}}_{i,k}}$
${{{\dot{\hat{x}}}}_{i,n}}={{u}_{i}}+\hat{W}_{i,n}^{T}{{E}_{i,n}}\left({{{\hat{\bar{x}}}}_{i,n}}\right)+{{q}_{i,n}}{{\psi}_{i,1}}+{{{\hat{\varpi}}}_{i,n}}$ (7)
with $k=1,...,n-1$.
Define the estimate of external perturbations as $\hat{\varpi}_{i}=[\hat{\varpi}_{i,1}^{T},...,\hat{\varpi}_{i,n}^{T}]^{T}$. and since the external perturbations are unknown, with the aim of achieving consensus control, a novel auxiliary variable is defined as
${{\tau}_{i,l}}={{\varpi}_{i,l}}-{{\kappa}_{i,l}}{{x}_{i,l}}$ (8)
where $\kappa_{i,l}$ is a designed positive parameter.
The perturbation observer is designed as
${{{\hat{\varpi}}}_{i,l}}$ $={{{\hat{\tau}}}_{i,l}}+{{\kappa}_{i,l}}{{{\hat{x}}}_{i,l}}$
${{{\dot{\hat{\tau}}}}_{i,l}}$ $=-{{\kappa}_{i,l}}(\hat{W}_{i,l}^{T}{{E}_{i,l}}({{{\hat{\bar{x}}}}_{i,l}})+{{{\hat{\tau}}}_{i,l}}+{{\kappa}_{i,l}}{{{\hat{x}}}_{i,l}}+{{{\hat{x}}}_{i,l+1}})$ (9)
where ${{x}_{i,n+1}}={{u}_{i}}$ and $\hat{W}_{i,l}$ represents the estimation of $W_{i,l}^{*}$.
II-C Graph Theory
This article examines MASs comprising $\mathcal{M}$ follower agents and one leaders. The interaction among these agents is represented using a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of nodes, $\mathcal{V}=\{1,...,N\}$, and $\mathcal{E}$ represents the set of edges, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. By the way, the set $\mathcal{E}=(\mathcal{V}_{i},\mathcal{V}_{j})\in\mathcal{E}$ means that agent $j$ can gain information from agent $i$. The adjacent matrix $\mathcal{A}=[a_{i,j}]\in R^{N\times N}$, where
$a_{i,j}=\left\{\begin{aligned} &1,\text{ }if\text{ }(\mathcal{V}_{i},\mathcal{V}_{j})\in\mathcal{E}\\ &0,\text{ }if\text{ }(\mathcal{V}_{i},\mathcal{V}_{j})\notin\mathcal{E}\end{aligned}\right.$ (10)
Define the Laplacian matrix $\mathcal{L}$ as $\mathcal{L}=\mathcal{D}-\mathcal{A}$, where $\mathcal{D}=diag({{d_{1}},...,{d_{N}}})$ is defined as the in-degree matrix of graph $\mathcal{G}$, ${d_{i}}=\sum\nolimits_{j=1,j\neq i}^{N}{{a_{ij}}}$. And clarify the diagonal matrix $\mathcal{B}=diag\left({{\mathcal{C}_{1}},...,{\mathcal{C}_{N}}}\right)$, if the leader 0 is capable of transmitting information to agent $i$, then ${\mathcal{C}_{i}}>0$. If not, ${\mathcal{C}_{i}}=0$ holds.
The previous section detailed the system setup process, while Fig. 1 presents the block diagram of the adaptive observer-based control framework with event-triggered strategies.
Figure 1: Block diagram of the adaptive observer-based control with event-triggered strategies.
$Lemma\text{ 1}$[17]: For any specified positive definite matrix ${{H}_{i}}=H_{i}^{T}>0$, with a symmetric positive matrix ${{F}_{i}}$, the equality $B_{i}^{T}{{F}_{i}}+{{F}_{i}}{{B}_{i}}=-2{{H}_{i}}$ holds.
$Lemma\text{ 2}$[18]: The inequality $0\leq\left|\chi\right|-\chi\tanh(\frac{\chi}{{{\chi}_{0}}})\leq 0.2785{{\chi}_{0}}$ always holds with any specified parameter $\chi\in R$ and ${{\chi}_{0}}>0$.
$Assumption\text{ }1$[19] : The unmeasured external perturbations are all bounded, i.e., the inequality $\left\|{{\varpi}_{i,j}}\right\|\leq\varpi_{i,j}^{0}$ holds, where $\varpi_{i,j}^{0}$ defines an positive parameter.
$Assumption\text{ }2$[20] : In this MASs, the leader’s desired reference signal is both measurable and smooth, with ${{y}_{r}}\left(t\right)$, ${{\dot{y}}_{r}}\left(t\right)$ being bounded.
III The Controller Design
Define the graph-based errors $z_{i,k}$ and the boundary layer errors $e_{i,k}$ for the ith follower as
$z_{i,1}$ $=\sum_{j=1}^{N}a_{i,j}(y_{i}-y_{j})+b_{i}(y_{i}-y_{r})$
$z_{i,k}$ $=x_{i,k}-\bar{\alpha}_{i,k}$ (11)
$e_{i,k}=\bar{\alpha}_{i,k}-\alpha_{i,k}$ (12)
where $k=2,...,n$. The terms $\alpha_{i,k}$ and $\bar{\alpha}_{i,k}$ respectively denote the virtual control and its filtered counterpart.
$Step\text{ }1$: The differential of $z_{i,1}$ is
${{\dot{z}}_{i,1}}=$ $\left({{d}_{i}}+{{b}_{i}}\right)\left({{{\hat{x}}}_{i,2}}+{{\psi}_{i,2}}\left(t\right)+{{\varpi}_{i,1}\left(t\right)+{{f}_{i,1}}\left({{{\bar{x}}}_{i,1}}\right)}\right)-{{b}_{i}}{{\dot{y}}_{r}}$
$-\sum\limits_{j=1}^{N}{{{a}_{i,j}}\left({{{\hat{x}}}_{j,2}}+{{\psi}_{j,2}}\left(t\right)+{{\varpi}_{j,1}}\left(t\right)+{{f}_{j,1}}\left({{{\bar{x}}}_{j,1}}\right)\right)}$
where $d_{i}=\sum_{j=1}^{N}a_{ij}$.
The Lyapunov function is determined as
${{V}_{i,1}}$ $=\frac{z_{i,1}^{2}}{2}+\frac{\tilde{\tau}_{i,1}^{2}}{2}+\frac{1}{2{{\eta}_{i,1}}}\tilde{W}_{i,1}^{T}{{\tilde{W}}_{i,1}}+\frac{1}{2{{o}_{i}}}\tilde{\Theta}_{i}^{2}+{{V}_{0}}$
$V_{0}$ $=\frac{1}{2}\psi_{i}^{T}(F_{i}\otimes I_{m})\psi_{i}$ (14)
where $\psi_{i}=[\psi_{i,1}^{T},...,\psi_{i,n}^{T}]^{T}$, $\eta_{i,1}>0$ and $o_{i}>0$ define designed parameters.
Define an auxiliary function as
${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)=$ $-\sum\limits_{j=1}^{N}{{{a}_{i,j}}}\left({{{\hat{x}}}_{j,2}}+{{f}_{j,1}}\left({{{\bar{x}}}_{j,1}}\right)+{{\varpi}_{j,1}}+{{\psi}_{j,2}}\right)$
$+\left({{d}_{i}}+{{b}_{i}}\right)\left({{f}_{i,1}}\left({{{\bar{x}}}_{i,1}}\right)+{{\varpi}_{i,1}}+{{\psi}_{i,2}}\right)-{{b}_{i}}{{\dot{y}}_{r}}.$ (15)
${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)$ is estimated by RBF NNs:
${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)=$ $K_{i,1}^{*T}{{E}_{i,1}}\left({{T}_{i,1}}\right)+{{\delta}_{i,1}}\left({{T}_{i,1}}\right)$
$\left|{{\delta}_{i,1}}\left({{T}_{i,1}}\right)\right|\leq$ $\bar{\delta}_{i,1}$ (16)
where $\bar{\delta}_{i,1}$ defines an unknown positive parameter.
Based on the Young’s inequality, the subsequent inequality can be derived:
${{z}_{i,1}}{{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)\leq$ $\frac{\Theta_{i}^{*}}{2c_{i,1}^{2}}z_{i,1}^{2}E_{i,1}^{T}\left({{T}_{i,1}}\right){{E}_{i,1}}\left({{T}_{i,1}}\right)$
$+\frac{c_{i,1}^{2}}{2}+\frac{z_{i,1}^{2}}{2}+\frac{\bar{\delta}_{i,1}^{2}}{2}.$ (17)
where ${c}_{i,1}$ is a positive designed parameter and ${{T}_{i,1}}={{[{{y}_{r}},{{\dot{y}}_{r}},\hat{x}_{i,1}^{T},\hat{x}_{j,1}^{T},{{\hat{\varpi}}_{i,1}}^{T},{{\hat{\varpi}}_{j,1}}^{T}]}^{T}}$ denotes the input of RBF NNs.
In accordance with the Young’s inequality and Assumption 1, one has
$-{{\kappa}_{i,1}}{{\tilde{\tau}}_{i,1}}{{\sigma}_{i,1}}\leq$ $\frac{1}{2}\tilde{\tau}_{i,1}^{2}+\frac{\kappa_{i,1}^{2}\sigma_{0,1}^{2}}{2}$ (18)
$-{{\kappa}_{i,1}}{{\tilde{\tau}}_{i,1}}{{\psi}_{i,2}}\leq$ $\frac{1}{2}\tilde{\tau}_{i,1}^{2}+{{{\kappa_{i,1}^{2}\left\|{{\psi}_{i}}\right\|}^{2}}}$ (19)
$-\kappa_{i,1}^{2}{{\tilde{\tau}}_{i,1}}{{\psi}_{i,1}}\leq$ $\frac{1}{2}\tilde{\tau}_{i,1}^{2}+{{{\kappa_{i,1}^{4}\left\|{{\psi}_{i}}\right\|}^{2}}}$ (20)
$\psi_{i}^{T}\left({{F}_{i}}\otimes{{I}_{m}}\right){{\varphi}_{i}}\leq$ $\frac{1}{2}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}{{\left\|\varphi_{i}^{0}\right\|}^{2}}+\frac{1}{2}{{\left\|{{\psi}_{i}}\right\|}^{2}}$ (21)
$\psi_{i}^{T}\left({{F}_{i}}\otimes{{I}_{m}}\right){{\tilde{\varpi}}_{i}}\leq$ $\frac{1}{2}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}{{\left\|\tilde{\varpi}_{i}^{0}\right\|}^{2}}+\frac{1}{2}{{\left\|{{\psi}_{i}}\right\|}^{2}}$ (22)
The design of the virtual control law and the adaptive law is formulated as
${{\alpha}_{i,2}}=$ $\frac{1}{{{d}_{i}}+{{b}_{i}}}({{r}_{i,1}}{{z}_{i,1}}-\frac{{{z}_{i,1}}}{2}-\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,1}^{2}}z_{i,1}E_{i,1}^{T}\left({{T}_{i,1}}\right){{E}_{i,1}}\left({{T}_{i,1}}\right))$ (23)
${{\dot{\hat{W}}}_{i,1}}=$ $-{{h}_{i,1}}{{\hat{W}}_{i,1}}-{{\eta}_{i,1}}{{\tilde{\tau}}_{i,1}}{{\kappa}_{i,1}}{{E}_{i,1}}\left({{{\hat{\bar{x}}}}_{i,1}}\right)$ (24)
where ${{h}_{i,1}}$ and ${{r}_{i,1}}$ are designed parameters.
Then, considering Lemma 1 and the above formulas, one has
$\dot{V}_{i,1}\leq$ ${{r}_{i,1}}z_{i,1}^{2}+\left({{d}_{i}}+{{b}_{i}}\right){{z}_{i,1}}\left({{z}_{i,2}}+{{e}_{i,2}}\right)+{{\tilde{\tau}}_{i,1}}{{\dot{\varpi}}_{i,1}}$
$+(\frac{3-2{{\kappa}_{i,1}}}{2})\tilde{\tau}_{i,1}^{2}+{{\iota}_{i,1}}+\frac{{{h}_{i,1}}}{{{\eta}_{i,1}}}\tilde{W}_{i,1}^{T}{{\hat{W}}_{i,1}}$
$+{{\tilde{\Theta}}_{i}}(\frac{1}{2c_{i,1}^{2}}z_{i,1}^{2}E_{i,1}^{T}({{T}_{i,1}}){{E}_{i,1}}({{T}_{i,1}})-\frac{{{{\dot{\hat{\Theta}}}}_{i}}}{{{o}_{i}}})$
$-\psi_{i}^{T}({{H}_{i}}\otimes{{I}_{m}}){{\psi}_{i}}+(1+\kappa_{i,1}^{2}+\kappa_{i,1}^{4}){{\left\|{{\psi}_{i}}\right\|}^{2}}$ (25)
where ${{\iota}_{i,1}}=\frac{c_{i,1}^{2}}{2}+\frac{\bar{\delta}_{i,1}^{2}}{2}+\frac{\kappa_{i,1}^{2}\sigma_{0,1}^{2}}{2}+\frac{1}{2}{{\left\|\varphi_{i}^{0}\right\|}^{2}}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}+\frac{1}{2}{{\left\|\tilde{\varpi}_{i}^{0}\right\|}^{2}}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}$.
Then, process ${{\alpha}_{i,2}}$ through the first-order low-pass filter, resulting in ${{\bar{\alpha}}_{i,2}}$ as follows:
${{\bar{\alpha}}_{i,2}}\left(0\right)={{\alpha}_{i,2}}\left(0\right)$
${{m}_{i,2}}{{\dot{\bar{\alpha}}}_{i,2}}+{{\bar{\alpha}}_{i,2}}={{\alpha}_{i,2}}$ (26)
where $m_{i,2}$ is a small positive designed parameter.
$Step\text{ }k$: Differentiating ${{z}_{i,k}}$, one has
${{\dot{z}}_{i,k}}=$ ${{z}_{i,k+1}}+{{s}_{i,k+1}}+\hat{W}_{i,k}^{T}{{E}_{i,k}}\left({{{\hat{\bar{x}}}}_{i,k}}\right)+{{q}_{i,k}}{{\psi}_{i,1}}$
$+{{\hat{\varpi}}_{i,k}}(t)+{{\alpha}_{i,k+1}}-{{\dot{\bar{\alpha}}}_{i,k}}$ (27)
where $k=2,...,n-1$.
Choose the Lyapunov function as
${{V}_{i,k}}=\frac{z_{i,k}^{2}}{2}+\frac{\tilde{\tau}_{i,k}^{2}}{2}+\frac{1}{2{{\eta}_{i,k}}}\tilde{W}_{i,k}^{T}{{\tilde{W}}_{i,k}}$ (28)
where ${{\eta}_{i,k}}$ is a positive parameter.
By utilizing RBF NNs, one has
${{\bar{Z}}_{i,k}}\left({{T}_{i,k}}\right)=$ $K_{i,k}^{*T}{{E}_{i,k}}\left({{T}_{i,k}}\right)+{{\delta}_{i,k}}\left({{T}_{i,k}}\right)$
$\left|{{\delta}_{i,k}}\left({{T}_{i,k}}\right)\right|\leq$ $\bar{\delta}_{i,k}$ (29)
where $\bar{\delta}_{i,k}$ denotes an unknown positive parameter.
The design of the virtual control law and the adaptive law is formulated as follows
${\alpha_{i,k+1}}=$ ${r_{i,k}}{z_{i,k}}-\frac{{{z_{i,k}}}}{2}+\frac{{{\alpha_{i,k}}-{{\bar{\alpha}}_{i,k}}}}{{{m_{i,k}}}}-{q_{i,k}}{\psi_{i,1}}$
$-\frac{{{{\hat{\Theta}}}_{i}}}{{2c_{i,k}^{2}}}{z_{i,k}}E_{i,k}^{T}\left({{T_{i,k}}}\right){E_{i,k}}\left({{T_{i,k}}}\right)$ (30)
${{\dot{\hat{W}}}_{i,k}}=$ $-{{h}_{i,k}}{{\hat{W}}_{i,k}}-{{\eta}_{i,k}}{{\tilde{\tau}}_{i,k}}{{\kappa}_{i,k}}{{E}_{i,k}}\left({{{\hat{\bar{x}}}}_{i,k}}\right)$ (31)
where ${{T}_{i,k}}={{[{{y}_{r}},{{\hat{\Theta}}_{i}},\hat{\bar{x}}_{i,k}^{T},\hat{\bar{x}}_{j,k}^{T},\hat{\varpi}_{i,k}^{T},\hat{\varpi}_{j,k}^{T}]}^{T}}$, ${{r}_{i,k}}$ and ${{h}_{i,k}}$ are designed parameters.
In terms of the Young’s inequality, one has
${{\dot{V}}_{i,k}}\leq$ ${{r}_{i,k}}z_{i,k}^{2}+{{\left(\kappa_{i,k}^{2}+\kappa_{i,k}^{4}\right)\left\|{{\psi}_{i}}\right\|}^{2}}+{{z}_{i,k}}\left({{z}_{i,k+1}}+{{e}_{i,k+1}}\right)$
$+\frac{{{h}_{i,k}}}{{{\eta}_{i,k}}}\tilde{W}_{i,k}^{T}{{\hat{W}}_{i,k}}+(\frac{3-2{{\kappa}_{i,k}}}{2})\tilde{\tau}_{i,k}^{2}+{{\tilde{\tau}}_{i,k}}{{\dot{\varpi}}_{i,k}}$
$+\frac{{{{\tilde{\Theta}}}_{i}}}{2c_{i,k}^{2}}z_{i,k}^{2}E_{i,k}^{T}\left({{T}_{i,k}}\right){{E}_{i,k}}\left({{T}_{i,k}}\right)+{{\iota}_{i,k}}$ (32)
where ${{\iota}_{i,k}}=\frac{c_{i,k}^{2}}{2}+\frac{\bar{\delta}_{i,k}^{2}}{2}+\frac{\kappa_{i,k}^{2}\sigma_{0,k}^{2}}{2}$, and $c_{i,k}$ denotes a positive parameter.
In line with Step 1, the first-order low-pass filter is implemented as follows:
${{\bar{\alpha}}_{i,k+1}}\left(0\right)={{\alpha}_{i,k+1}}\left(0\right)$
${{m}_{i,k+1}}{{\dot{\bar{\alpha}}}_{i,k+1}}+{{\bar{\alpha}}_{i,k+1}}={{\alpha}_{i,k+1}}.$ (33)
where ${{m}_{i,k+1}}$ denotes a slight positive designed parameter.
$Step\text{ }n$: Differentiating ${{z}_{i,n}}$, one gets
${{\dot{z}}_{i,n}}=$ $u_{i}+\hat{W}_{i,n}^{T}{{E}_{i,n}}\left({{{\hat{\bar{x}}}}_{i,n}}\right)+{{q}_{i,n}}{{\psi}_{i,1}}+{{\hat{\varpi}}_{i,n}}(t)-{{\dot{\bar{\alpha}}}_{i,n}}$ (34)
Choose the Lyapunov function as
${{V}_{i,n}}=\frac{z_{i,n}^{2}}{2}+\frac{\tilde{\tau}_{i,n}^{2}}{2}+\frac{1}{2{{\eta}_{i,n}}}\tilde{W}_{i,n}^{T}{{\tilde{W}}_{i,n}}$ (35)
where ${{\eta}_{i,n}}$ denotes a designed positive parameter.
By utilizing RBF NNs, one has
${{\bar{Z}}_{i,n}}\left({{T}_{i,n}}\right)=$ $K_{i,n}^{*T}{{E}_{i,n}}\left({{T}_{i,n}}\right)+{{\delta}_{i,n}}\left({{T}_{i,n}}\right)$
$\left|{{\delta}_{i,n}}\left({{T}_{i,n}}\right)\right|\leq$ $\bar{\delta}_{i,n}$ (36)
where $\bar{\delta}_{i,n}$ denotes an unknown positive parameter.
The control law and adaptive parameters are constructed as
${\alpha_{i,n+1}}=$ $r_{i,n}{z_{i,n}}-\frac{{{z_{i,n}}}}{2}+\frac{{{\alpha_{i,n}}-{{\bar{\alpha}}_{i,n}}}}{{{m_{i,n}}}}-{q_{i,n}}{\psi_{i,1}}$
$-\frac{{{{\hat{\Theta}}}_{i}}}{{2c_{i,n}^{2}}}{z_{i,n}}E_{i,n}^{T}\left({{T_{i,n}}}\right){E_{i,n}}\left({{T_{i,n}}}\right)$ (37)
${{\dot{\hat{W}}}_{i,n}}=$ $-{{h}_{i,n}}{{\hat{W}}_{i,n}}-{{\eta}_{i,n}}{{\tilde{\tau}}_{i,n}}{{\kappa}_{i,n}}{{E}_{i,n}}\left({{{\hat{\bar{x}}}}_{i,n}}\right)$ (38)
${{\dot{\hat{\Theta}}}_{i}}=$ $-{{{\lambda}}_{i}}{{\hat{\Theta}}_{i}}+\sum\limits_{k=1}^{n}{\frac{{{o}_{i}}}{2c_{i,k}^{2}}z_{i,k}^{2}E_{i,k}^{T}\left({{T}_{i,k}}\right){{E}_{i,k}}\left({{T}_{i,k}}\right)}$ (39)
where ${{r}_{i,n}}$, ${{h}_{i,n}}$ and ${{\lambda}}_{i}$ are designed parameters.
III-A Fixed-threshold strategy
In this strategy, the adaptive event-triggered controller is reformulated as
${{w}_{i}}\left(t\right)={{\alpha}_{i,n+1}}-{{\bar{\pi}}_{i}}\tanh(\frac{{{z}_{i,n}}{{{\bar{\pi}}}_{i}}}{{{\mu}_{i}}})$ (40)
The triggering condition is defined as
${{u}_{i}}\left(t\right)=$ $w_{i}\left({{t}_{s}}\right)\text{ },\text{ }\forall\text{t}\in\left[{{t}_{s}},{{t}_{s+1}}\right)$ (41)
${{t}_{s+1}}=$ $\inf\left\{\text{t}\in R||{{\vartheta}_{i}}\left(t\right)|\geq{{\pi}_{i}}\right\}\text{, }{{\text{t}}_{1}}=0$ (42)
where $\vartheta_{i}(t)=w_{i}(t)-u_{i}(t)$ represents the measurement-triggered error, while $t_{s}$ indicates the controller’s update time, with $s\in Z$. The parameters $\bar{\pi}_{i}$, $\pi_{i}$ and $\mu_{i}$ are designed and positive, $\bar{\pi}>\pi_{i}$. The control signal ${{u}_{i}}\left({{t}_{s+1}}\right)$ is applied to the system when condition (42) is activated. During the interval $\text{t}\in\left[{{t}_{s}},{{t}_{s+1}}\right)$, i.e. $\left|{{w}_{i}}\left(t\right)-{{u}_{i}}\left(t\right)\right|<{{\pi}_{i}}$, the controller maintains a constant value of ${{w}_{i}}\left({{t}_{s}}\right)$. For a function ${{\epsilon}_{i}}{{\left(t\right)}}$ that changes continuously over time that satisfies ${{\epsilon}_{i}}\left({{t}_{s}}\right)=0$ and ${{\epsilon}_{i}}\left({{t}_{s+1}}\right)=\pm 1$ with $\left|{{\epsilon}_{i}}\left(t\right)\right|\leq 1$, it follows that ${{w}_{i}}\left(t\right)={{u}_{i}}\left(t\right)+{{\epsilon}_{i}}\left(t\right){{\pi}_{i}}$.
Based on Lemma 2, the inequality $-{{\epsilon}_{i}}(t){{\pi}_{i}}{{z}_{i,n}}-{{\bar{\pi}}_{i}}{{z}_{i,n}}\tanh(\frac{{{z}_{i,n}}{{{\bar{\pi}}}_{i}}}{{{\mu}_{i}}})\leq 0.2875{{\mu}_{i}}$ holds, then, differentiate ${{V}_{i,n}}$ by utilizing the Young’s inequality and (36)-(38), one gets
$\dot{V}_{i,n}\leq$ ${{r}_{i,n}}z_{i,n}^{2}+(\frac{3-2{{\kappa}_{i,n}}}{2})\tilde{\tau}_{i,n}^{2}+{{\tilde{\tau}}_{i,n}}{{\dot{\varpi}}_{i,n}}$
$+\frac{{{{\tilde{\Theta}}}_{i}}}{2c_{i,n}^{2}}z_{i,n}^{2}E_{i,n}^{T}\left({{T}_{i,n}}\right){{E}_{i,n}}\left({{T}_{i,n}}\right)+{{\iota}_{i,n}}$
$+\frac{{{h}_{i,n}}}{{{\eta}_{i,n}}}\tilde{W}_{i,n}^{T}{{\hat{W}}_{i,n}}+{{\left(\kappa_{i,n}^{2}+\kappa_{i,n}^{4}\right)\left\|{{\psi}_{i}}\right\|}^{2}}$ (43)
where ${{\iota}_{i,n}}=\frac{c_{i,n}^{2}}{2}+\frac{\bar{\delta}_{i,n}^{2}}{2}+\frac{\kappa_{i,n}^{2}\sigma_{0,n}^{2}}{2}+0.2785{{\mu}_{i}}$, $c_{i,n}$ is a positive parameter.
III-B Relative-threshold strategy
In the fixed-threshold strategy, it is observed that the threshold $\pi_{i}$ remains constant regardless of the control signal’s magnitude. Nevertheless, when addressing a stabilization problem, adopting a variable threshold for the triggering condition becomes advantageous, according to the approach in [21][22]. Specifically, when the control signal $u_{i}$ is huge, a greater error of measurement can be tolerated, allowing for longer update intervals. Conversely, as the system states stabilize towards equilibrium, with $u_{i}$ approaching zero, a smaller threshold enables more accurate control, enhancing overall system performance. Then, it is proposed the following adaptive event-triggered controller:
${{w}_{i}}\left(t\right)=$ $-(1+\Delta_{i})({{\alpha}_{i,n+1}}tanh(\frac{z_{i,n}\alpha_{i,n+1}}{\mu_{i}})$
$+{{\bar{\pi}}^{*}_{i}}\tanh(\frac{{{z}_{i,n}}{{{\bar{\pi}}}^{*}_{i}}}{{{\mu}_{i}}}))$ (44)
The triggering condition is defined as
${{u}_{i}}\left(t\right)=$ $w_{i}\left({{t}_{s}}\right)\text{ },\text{ }\forall\text{t}\in\left[{{t}_{s}},{{t}_{s+1}}\right)$ (45)
${{t}_{s+1}}=$ $\inf\left\{\text{t}\in R||{{\vartheta}_{i}}\left(t\right)|\geq\Delta_{i}|u_{i}(t)|+{{\pi}^{*}_{i}}\right\}$ (46)
where $t_{s}$ represents the time when the controller is updated, $s\in Z$, $\mu_{i}$, $\Delta_{i}$, $10<\Delta_{i}<$ 1, $\pi^{*}_{i}>0$, and $\bar{\pi}^{*}_{i}>\pi^{*}_{i}/(1-\Delta_{i})$ are all positive designed parameters.
From (49), we have $w_{i}(t)=(1+\rho_{1,i}(t))u_{i}(t)+\rho_{2,i}(t)\pi_{i}^{*}$ in the interval $[t_{s},t_{s+1}]$, where $\rho_{1,i}(t)$ and $\rho_{2,i}(t)$ denote time-varying parameters satisfying $|\rho_{1,i}(t)|\leq 1$ and $|\rho_{2,i}(t)|\leq 1$.
Thus, one has
${{u}_{i}}(t)=\frac{w_{i}(t)}{1+\rho_{1,i}(t)\Delta_{i}}-\frac{\rho_{2,i}(t)\pi_{i}^{*}}{1+\rho_{1,i}(t)\Delta_{i}}$ (47)
Since $\forall a_{i}\in R$ and $\mu_{i}>0$, from (45) we get $z_{i,n}w_{i}(t)\leq 0$, one has
$\frac{z_{i,n}w_{i}(t)}{1+\rho_{1,i}(t)\Delta_{i}}\leq\frac{z_{i,n}w_{i}(t)}{1+\Delta_{i}}$ (48)
$\frac{\rho_{2,i}\pi_{i}^{*}}{1+\rho_{1,i}(t)\Delta_{i}}\leq\frac{\pi_{i}^{*}}{1-\Delta_{i}}$ (49)
According to the above analysis, similar to (43), it can be obtained the same derivative result of $\dot{V}_{i,n}$ with different ${{\iota}_{i,n}}$, where ${{\iota}_{i,n}}=\frac{c_{i,n}^{2}}{2}+\frac{\bar{\delta}_{i,n}^{2}}{2}+\frac{\kappa_{i,n}^{2}\sigma_{0,n}^{2}}{2}+0.557{{\mu}_{i}}$.
III-C Switch-threshold strategy
In this portion, it is introduced a switch-threshold strategy. The relative-threshold approach adjusts the threshold based on the control signal’s magnitude, allowing for longer update intervals when the signal is huge and more precise control as the signal nears zero, thus improving performance. However, a very large control signal can lead to significant measurement errors and abrupt changes during updates. In contrast, the fixed-threshold strategy maintains a consistent upper limit on measurement errors, regardless of signal size. Considering these aspects, we propose the following switched-threshold strategy. The triggering condition is defined as follows:
${{u}_{i}}\left(t\right)=w_{i}\left({{t}_{s}}\right)\text{ },\text{ }\forall\text{t}\in\left[{{t}_{s}},{{t}_{s+1}}\right)$ (50)
${t}_{s+1}=\left\{\begin{aligned} &\inf\left\{\text{t}\in R||{{\vartheta}_{i}}\left(t\right)|\geq\Delta_{i}|u_{i}(t)|+{{\pi}^{*}_{i}}\right\},|u_{i}|\geq G\\ &\inf\left\{\text{t}\in R||{{\vartheta}_{i}}\left(t\right)|\geq{{\pi}_{i}}\right\},|u_{i}|<G\end{aligned}\right.$ (51)
where $G$ is the designed switching gate.
For this switch-threshold strategy, one has
$\bar{\vartheta}_{i}=sup|\vartheta_{i}(t)|\leq max\{(\Delta_{i}|u_{i}(t)|+{{\pi}^{*}_{i}}),{{\pi}_{i}}\}$ (52)
where $t\in[t_{s},t_{s+1})$.
Since the switch-threshold strategy employs an identical control law to both the fixed-threshold and relative-threshold strategies, the ultimate bounds for tracking and stabilization errors remain consistent with those of the two previously mentioned strategies.
IV Stability Analysis
Considering the previous analysis, we have developed observers and designed an adaptive consensus tracking controller via RBF NNs, backstepping method and three event-triggered strategies. In this section, we demonstrate the boundedness of all signals by employing Lyapunov stability theory. In addition, the event-triggered strategies are proven to be effective, and the Zeno behavior is avoided.
$Theorem\text{ 1}$: For the nonlinear MASs in equation (1) with the event-triggered controllers in equation (40)(44), if the initial condition satisfies $V(0)\leq\Omega$, the error signals $z_{i,k}$, $\tilde{c}_{i,k}$, ${{\tilde{W}}_{i,k}}$, ${{\tilde{\Theta}}_{i}}$, ${{e}_{i,k}}$, and ${{\psi}_{i,k}}$ are all uniformly bounded. Furthermore, the errors of consensus tracking between the followers’ outputs and the leader’s trajectory signal can be minimized to a specified range.
$Proof$: See Appendix A.
V Illustrative Example
This section will examine an illustrative example to validate the performance and effectiveness of the proposed approach. The system comprises one leader agent and four follower agents. The dynamics of each agent are described by:
${{{\dot{x}}}_{i,1}}={{x}_{i,2}}+{{f}_{i,1}}\left({{{\bar{x}}}_{i,1}}\right)+{{\xi}_{i,1}}$
${{{\dot{x}}}_{i,2}}={{u}_{i}}+{{f}_{i,2}}\left({{{\bar{x}}}_{i,2}}\right)+{{\xi}_{i,2}}$
${{y}_{i}}(t)={{x}_{i,1}}$ (53)
where ${{f}_{i,1}}\left({{{\bar{x}}}_{i,1}}\right)=0.8{{x}_{i,1}}{{e}^{-1.4x_{i,2}^{2}}}$, ${{f}_{i,2}}\left({{{\bar{x}}}_{i,2}}\right)=-0.5x_{i,1}^{2}\cos\left({{x}_{i,2}}\right)$, ${{\xi}_{i,1}}=0.8{{x}_{i,1}}\sin\left({{x}_{i,2}}\right)\cos^{2}\left(t\right)$ and ${{\xi}_{i,2}}=0.2{{x}_{i,2}}\cos\left({{x}_{i,1}}\right)\cos^{2}\left(t\right)$, $i=1,...,4$. The leader’s trajectory signal is defined as ${{y}_{r}}=-0.5\sin(4t)\cos(2t)$.
The communication topology is illustrated in Fig. 1. The connection matrix linking the leader to the followers is represented as $\mathcal{B}=diag\{0,1,0,0\}$. And the adjacency matrix $\mathcal{A}$ and the Laplacian matrix $\mathcal{L}$ are presented as follows:
$\mathcal{A}=\left[\begin{array}[]{cccc}0&1&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 1&0&0&0\\ \end{array}\right]$, $\mathcal{L}=\left[\begin{array}[]{cccc}1&-1&0&0\\ 0&0&0&0\\ 0&-1&1&0\\ -1&0&0&1\\ \end{array}\right]$.
The initial conditions of the four followers and state observers are selected as follows:
${{{\bar{x}}}_{1,2}}(0)={{[0.2,\text{ }0]}^{T}}$ ${{{\hat{\bar{x}}}}_{1,2}}(0)={{[0.3,\text{ 1.7}]}^{T}}$
${{{\bar{x}}}_{2,2}}(0)={{[-0.2,\text{ 0}]}^{T}}$ ${{{\hat{\bar{x}}}}_{2,2}}(0)={{[-0.5,\text{ 1.7}]}^{T}}$
${{{\bar{x}}}_{3,2}}(0)={{[0.1,\text{ 0}]}^{T}}$ ${{{\hat{\bar{x}}}}_{3,2}}(0)={{[0,\text{ -4}]}^{T}}$
${{{\bar{x}}}_{4,2}}(0)={{[-0.3,\text{ 0}]}^{T}}$ ${{{\hat{\bar{x}}}}_{4,2}}(0)={{[0,\text{ -4}]}^{T}}$
Based on the parameter selection guidelines outlined in the stability analysis, the quantitative values of the design parameters are as follows: for fix-threshold strategy, ${{\pi}_{i}}=2.5$, ${{\bar{\pi}}_{i}}=4$, ${{\mu}_{i}}=5.4$, for relative-threshold strategy, ${{\pi}_{i}^{*}}=2$, ${{\bar{\pi}}_{i}^{*}}=4$, $\Delta_{i}=0.245$, for switch-threshold strategy $G=6$, for the first order low pass filter, $m_{i}=0.005$, for other parameters, ${{h}_{i,1}}={{h}_{i,2}}=50$, ${{r}_{i,1}}={{r}_{i,2}}=-100$, ${{c}_{i,1}}={{c}_{i,2}}=100$, ${{\eta}_{i,1}}={{\eta}_{i,2}}=0.01$, ${{q}_{i,1}}=350$, ${{q}_{i,2}}=0.5$, ${{\lambda}_{i}}=120$ and ${{o}_{i}}=25$.
In Fig. 2, shortly after the process begins, all followers within the system consistently follow the leader, successfully achieving the adaptive consensus control objective. Fig. 3 is presented to illustrate that the tracking error has converged to zero. The total triggering number for sample-data strategy is 5,000 times, and the triggering numbers for the three threshold strategies are detailed in Table I. Figs. 4 shows that the interval of events respectively triggered by three strategies. According to the statistics, it is clear that under the fixed-threshold strategy, the number of triggers is the highest. Under the relative-threshold strategy, the number of triggers is the lowest, while under the switch-threshold strategy, the number of triggers falls and both types of triggers play an indispensable role in this strategy.
Table I: Comparison of three event-triggered strategies.
Fixed-threshold Switch-threshold Relative-threshold
Follower 1 $364$ $344(255+89)$ $310$
Follower 2 $296$ $281(220+61)$ $277$
Follower 3 $358$ $342(255+87)$ $308$
Follower 4 $453$ $420(306+114)$ $380$
Figure 2: Consensus tracking performance.
Figure 3: Trajectories of tracking errors $z_{i,1}$($i=1,2,3,4.$)
Fixed-threshold strategy
Switch-threshold strategy
Relative-threshold strategy
Figure 4: Event-triggered release instants and interval: three strategies.
$Remark\text{ }1$: The theoretical analysis and simulation results indicate that the fixed-threshold strategy maintains bounded measurement error regardless of control amplitude, reducing the risk of excessive control signals that could impair system performance. In contrast, the relative-threshold strategy adjusts controller update intervals based on control amplitude variations, minimizing control actions while maintaining tracking performance, thereby extending controller lifespan and improving resource utilization. The switch-threshold strategy combines the benefits of both approaches by regulating trigger occurrences at an intermediate level, adjusting activation frequency through a designed switching gate.
VI Conclusion
This article addresses the observer-based consensus tracking issue for nonlinear MASs via three event-triggered control strategies. The RBF NNs are utilized to estimate unidentified nonlinear functions, while observers handle unmeasurable states and unknown perturbations. Complexity is managed by incorporating a filter at each design step. Furthermore, by comprising three event-triggered strategies, Besides, three distinct strategies for controller updates are presented. Simulation results indicate that the proposed Fixed-threshold strategy can better ensure system performance while reducing the number of triggering instances. Relative-threshold strategy minimizes the number of triggering instances to conserve resources. Switch-threshold strategy effectively balances system performance with resource utilization for each agent. Future work will focus on the optimized event-triggered control for stochastic nonlinear MASs.
Acknowledgment
The authors wish to convey their appreciation to Ning Pang at Shanghai Jiao Tong University for insightful discussions about some of the issues addressed in this paper. This work is partially supported by the National Natural Science Foundation of China under grants 62303389 and 62373289, Guangdong Basic and Applied Basic Research Funding under grants 2022A151511076 and 2024A1515012586, Guangzhou Basic and Applied Basic Research Scheme under grants 2023A04J1067, together with Guangzhou-HKUST(GZ) Joint Funding Program under grants 2023A03J0678, 2023A03J0011, 2024A03J0618 and 2024A03J0680.
[PROOF OF THEOREM 1] Define the following $i=1,...,N$ and the overall Lyapunov function for the MASs (1) as
$V=\sum\limits_{i=1}^{N}{\sum\limits_{k=1}^{n}{{V_{i,k}}}}+\sum\limits_{i=1}^{N}\sum\limits_{k=1}^{n-1}{\frac{{e_{i,k+1}^{2}}}{2}}$ (54)
Based on the Young’s inequality, one gets
${{z}_{i,1}}\left({{z}_{i,2}}+{{e}_{i,2}}\right)\leq$ $z_{i,1}^{2}+\frac{1}{2}z_{i,2}^{2}+\frac{1}{2}e_{i,2}^{2}$ (55)
$\tilde{E}_{i,k}^{T}{{\hat{E}}_{i,k}}\leq$ $*\displaystyle-\frac{1}{2}\tilde{E}_{i,k}^{T}{{\tilde{E}}_{i,k}}+\frac{1}{2}E_{i,k}^{\text{*}T}E_{i,k}^{\text{*}}$ (56)
${{\tilde{\Theta}}_{i}}{{\hat{\Theta}}_{i}}\leq$ $\frac{1}{2}{{\left(\Theta_{i}^{*}\right)}^{2}}-\frac{1}{2}\tilde{\Theta}_{i}^{2}.$ (57)
Define the parameters $r_{i,1}$, $r_{i,2}$, $r_{i,k}$ and $r_{i,n}$ as ${{r}_{i,1}}=-\left({{d}_{i}}+{{b}_{i}}\right)+r_{i,1}^{*}$, ${{r}_{i,2}}=-\frac{{{d}_{i}}+{{b}_{i}}}{2}-1+r_{i,2}^{*}$, ${{r}_{i,k}}=-\frac{3}{2}+r_{i,k}^{*}$ and ${{r}_{i,n}}=-\frac{1}{2}+r_{i,n}^{*}$ with $k=3,...,n-1$.
While $\left\|{{{\dot{\varpi}}}_{i,k}}\right\|\leq\kappa_{i,k}^{*}\left\|{{{\tilde{\tau}}}_{i,k}}\right\|$, one has
$\dot{V}\leq$ $\sum\limits_{i=1}^{N}\Bigg{\{}{\sum\limits_{k=1}^{n}{r_{i,k}^{*}}}z_{i,k}^{2}+{\sum\limits_{k=2}^{n-1}{\frac{e_{i,k+1}^{2}}{2}}}+\frac{{{d}_{i}}+{{b}_{i}}}{2}e_{i,2}^{2}$
$+{\sum\limits_{k=1}^{n-1}{{{e}_{i,k+1}}}}{{{\dot{e}}}_{i,k+1}}+2{\sum\limits_{k=1}^{n}{\kappa_{i,k}^{*}}}\tilde{\tau}_{i,k}^{2}+\frac{{{\lambda}_{i}}}{{2{o}_{i}}}({\Theta_{i}^{*}}^{2}-\tilde{\Theta}_{i}^{2})$
$+{\sum\limits_{k=1}^{n}{\frac{{{h}_{i,k}}}{{2{\eta}_{i,k}}}(W_{i,k}^{\text{*}T}W_{i,k}^{\text{*}}}}-\tilde{W}_{i,k}^{T}{{{\tilde{W}}}_{i,k}})+{\sum\limits_{k=1}^{n}{{{\iota}_{i,k}}}}$
$-\psi_{i}^{T}(({{H}_{i}}-(1+\sum\limits_{k=1}^{n}{(\kappa_{i,k}^{2}+\kappa_{i,k}^{4})}){{I}_{n}})\otimes{{I}_{m}}){{\psi}_{i}}\Bigg{\}}$ (58)
where $\kappa_{i,k}^{*}=(3-2{{\kappa}_{i,k}})/2$, $r_{i,k}^{*}$ and $\kappa_{i,k}^{*}$ denote unmeasured parameters for stability analysis, with $k=1,...,n$.
$\hbar(\bullet)$ is defined as the eigenvalue of the given matrix. Select the matrix ${{H}_{i}}$ such that ${{\hbar}_{\min}}[{{H}_{i}}-(1+\sum\limits_{k=1}^{n}{(\kappa_{i,k}^{2}+\kappa_{i,k}^{4})}){{I}_{n}}]={{\wp}_{i}}/2{{\hbar}_{\max}}({{F}_{i}})$, where ${{\wp}_{i}}$ denotes a positive parameter.
Differentiate with respect to time of ${{e}_{i,2}}$ and ${{e}_{i,k+1}}$, one has
${{\dot{e}}_{i,2}}=$ $-\frac{{{e}_{i,2}}}{{{m}_{i,2}}}+{\Gamma_{i,2}}$
${{\dot{e}}_{i,k+1}}=$ $-\frac{{{e}_{i,k+1}}}{{{m}_{i,k+1}}}+{\Gamma_{i,k+1}}$ (59)
where ${\Gamma_{i,2}}$ and ${\Gamma_{i,k+1}}$ can be represented as ${{\Gamma}_{i,2}}=-\frac{{{r}_{i,1}}{{{\dot{z}}}_{i,1}}}{{{d}_{i}}+{{b}_{i}}}+\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,1}^{2}\left({{d}_{i}}+{{b}_{i}}\right)}{{\dot{z}}_{i,1}}E_{i,1}^{T}\left({{T}_{i,1}}\right){{E}_{i,1}}\left({{T}_{i,1}}\right)+\frac{{{{\dot{z}}}_{i,1}}}{2\left({{d}_{i}}+{{b}_{i}}\right)}+\frac{{{{\hat{\Theta}}}_{i}}}{c_{i,1}^{2}\left({{d}_{i}}+{{b}_{i}}\right)}{{z}_{i,1}}E_{i,1}^{T}\left({{T}_{i,1}}\right){{\dot{E}}_{i,1}}\left({{T}_{i,1}}\right)$ and ${{\Gamma}_{i,k+1}}=\frac{{{\dot{e}}_{i,k}}}{{{m}_{i,k}}}+{{q}_{i,k}}{{\dot{\psi}}_{i,1}}+\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,k}^{2}}{{\dot{z}}_{i,k}}E_{i,k}^{T}\left({{T}_{i,k}}\right){{E}_{i,k}}\left({{T}_{i,k}}\right)-{{r}_{i,k}}{{\dot{z}}_{i,k}}+\frac{{{{\dot{z}}}_{i,k}}}{2}+\frac{{{{\hat{\Theta}}}_{i}}}{c_{i,k}^{2}}{{z}_{i,k}}E_{i,k}^{T}\left({{T}_{i,k}}\right){{\dot{E}}_{i,k}}\left({{T}_{i,k}}\right)$.
Inspired by the work presented in [20], we analyze the sets
${{\mathbb{M}}_{i,k}}=$$\Bigg{\{}\sum\limits_{i=1}^{N}{\sum\limits_{k=1}^{n}{(z_{i,k}^{2}+\tilde{\tau}_{i,k}^{2}+\frac{1}{{{\eta}_{i,k}}}\tilde{W}_{i,k}^{T}{{{\tilde{W}}}_{i,k}})}}+\sum\limits_{i=1}^{N}{\frac{1}{{{o}_{i}}}\tilde{\Theta}_{i}^{2}}$
$+\sum\limits_{i=1}^{N}{\psi_{i}^{T}({{F}_{i}}\otimes{{I}_{m}}){{\psi}_{i}}}+\sum\limits_{i=1}^{N}{\sum\limits_{k=1}^{n-1}{e_{i,k+1}^{2}}}\leq 2\Omega\Bigg{\}}$ (60)
where ${{\mathbb{M}}_{i,k}}$ is compact in ${{R}^{\dim\left({{\mathbb{C}}_{i,k}}\right)}}$, there exists an inequality $\left\|{{\Gamma}_{i,k+1}}\right\|\leq{{L}_{i,k+1}}$, where $L_{i,k+1}$ denotes a positive parameter.
Applying the Young’s inequality, one can get ${{\Gamma}_{i,k+1}}{{e}_{i,k+1}}\leq\frac{\Xi}{2}+\frac{L_{i,k+1}^{2}e_{i,k+1}^{2}}{2\Xi}$, where $\Xi$ is a positive parameter, with $k=1,...,n-1$.
Select the parameters as $\frac{1}{{{m}_{i,2}}}=\frac{{{d}_{i}}+{{b}_{i}}}{2}+\frac{L_{i,2}^{2}}{2\Xi}+m_{i,2}^{*}$ and $\frac{1}{{{m}_{i,k+1}}}=\frac{1}{2}+\frac{L_{i,k+1}^{2}}{2\Xi}+m_{i,k+1}^{*}$, where $m_{i,k+1}^{*}$ denotes an unknown positive parameter, with $k=2,...,n-1$.
With $l=2,...,n$ and $k=1,...,n$, define
$\beta=$ $\min\left\{-2r_{i,k}^{*},\text{ -4}\kappa_{i,k}^{*},\text{ }{{h}_{i,k}},\text{ }{{\wp}_{i}},\text{ }{{\lambda}_{i}},\text{ 2}{{m}_{i,l}^{*}}\right\}$ (61)
$\gamma=$ $\sum\limits_{i=1}^{N}{\sum\limits_{k=1}^{n}{({{\iota}_{i,k}+{\frac{{{h}_{i,k}}}{2{{\eta}_{i,k}}}W_{i,k}^{\text{*}T}W_{i,k}^{\text{*}}})}}}+\sum\limits_{i=1}^{N}({\sum\limits_{k=1}^{n-1}{\frac{\Xi}{2}+{\frac{{{\lambda}_{i}}}{2{{o}_{i}}}{{\left(\Theta_{i}^{*}\right)}^{2}}}}})$ (62)
To ensure the stability of the entire MASs, it is crucial to appropriately select parameter values while adhering to the following conditions: $-r_{i,k}^{*}>0,\text{ }-\kappa_{i,k}^{*}>0,\text{ }{{h}_{i,k}}>0,\text{ }{{\wp}_{i}}>0,\text{ }{{\lambda}_{i}}>0$ and ${{m}_{i,l}^{*}}>0.$
Then, one has
$\dot{V}\leq-\beta V+\gamma$ (63)
When set $\beta>\gamma/\Delta$, we get $\dot{V}<0$ on $V=\Delta$. Further, if at time $t=0$ the condition $V\leq\Delta$ holds, it follows that $V\leq\Delta$ for entire $t>0$. This demonstrates that the error signals ${{z}_{i,k}}$, ${{e}_{i,k}}$, ${{\tilde{\tau}}_{i,k}}$, ${{\tilde{W}}_{i,k}}$, ${{\tilde{\Theta}}_{i}}$ and ${{\psi}_{i,k}}$ are uniformly bounded. It is straightforward to derive the following:
$\frac{1}{2}{{\left\|{{\Upsilon}_{1}}\right\|}^{2}}\leq V\left(t\right)\leq{{e}^{-\beta t}}V\left(0\right)+\frac{\gamma}{\beta}\left(1-{{e}^{-\beta t}}\right)$ (64)
where ${{\Upsilon}}={{\left[\Upsilon_{1}^{T},\Upsilon_{2}^{T},\ldots,\Upsilon_{M}^{T}\right]}^{T}}$. Then, we have ${{\left\|{{\Upsilon}}\right\|}^{2}}\leq 2{{e}^{-\beta t}}V\left(0\right)+\frac{2\gamma}{\beta}\left(1-{{e}^{-\beta t}}\right)$.
Consequently, as time progresses, all consensus tracking errors will converge to a compact set defined as =$\Im=\{{{\Upsilon}_{1}}|\left\|{{\Upsilon}_{1}}\right\|\leq\sqrt{2\gamma/\beta}\}$. This means that the tracking errors among agents can be modified and reduced to an arbitrarily small range by increasing the parameter $\beta$. From (43), one has
${{\dot{\alpha}}_{i}}\left(t\right)={{\dot{w}}_{i,n+1}}-\frac{\bar{\pi}_{i}{{{\dot{z}}}_{i,n}}}{{{\cosh}^{2}}\left(\frac{{{z}_{i,n}}{{{\bar{\pi}}}_{i}}}{{{\mu}_{i}}}\right)}.$ (65)
According to [23], one gets
$\frac{d}{dt}\left|{{\vartheta}_{i}}\left(t\right)\right|$ $=\frac{d}{dt}{{\left({{\vartheta}_{i}}\left(t\right)\times{{\vartheta}_{i}}\left(t\right)\right)}^{\frac{1}{2}}}$
$=sign\left({{\vartheta}_{i}}\left(t\right)\right){{\dot{\vartheta}}_{i}}\left(t\right)\leq\left|{{{\dot{w}}}_{i}}\left(t\right)\right|.$ (66)
Based on the stability analysis, it is imperative that there exists a positive parameter $\varPi$ such that $|{{{\dot{w}}}_{i}}(t)|\leq\varPi$. Based on (41)(42), we have ${{\vartheta}_{i}}({{t}_{s}})=0$ and $\underset{{}}{\mathop{{{\lim}_{t\to{{t}_{s+1}}}}}}\,{{\vartheta}_{i}}(t)={{\pi}_{i}}$. Additionally, for the time interval $t\in[{{t}_{s}},{{t}_{s+1}})$, the lower bound for the inter-execution time is given by ${{t}^{*}}\geq{{\pi}_{i}}/{\varPi}$.
Following the same analysis in the proof of fixed-threshold strategy, according to (45)(46), the relative-threshold and switch-threshold strategy satisfy ${{t}^{*}}\geq({\Delta_{i}|u_{i}(t)|+{\pi}_{i}^{*}})/{\varPi}$ and ${{t}^{*}}\geq{max\{{\pi}_{i},{\pi}_{i}^{*}\}}/{\varPi}$, respectively. Consequently, the Zeno behaviour is proficiently avoided.